2. Use the following table to determine the absolute advantage and comparative advantage of China and the United States in producing wheat and textiles. Explain your answers. 
Answer #1

ANSWER:

United states has absolute advantage in both textiles and wheat. ($165,000 > $27,000 in textiles) and (12,260 > 300 in wheat)

opportunity cost of 1 wheat for united states = 165,000 / 12,260 = 13.45

opportunity cost of 1 wheat for china = 27,000 / 300 = 90

opportunity cost of 1 textile for united states = 12,260 / 165,000 = 0.074

opportunity cost of 1 textile for china = 300 / 27,000 = 0.011

so china has comparative advantage in producing textiles(as opportunity cost of producing 1 textile is less than that for united states that is 0.011 < 0.074) and so united states has comparative advantage in producing wheat.(lower opportunity cost in producing wheat. than china. that is 13.45 < 90)